OpenStudy (anonymous):

how do i find the sum of the infinite series ((-2)^n)/(e^(n+1))?

OpenStudy (anonymous):

\[\sum_{n =1}^{\infty} \frac{ (-2)^{n} }{ e^{n+1} }\]

geerky42 (geerky42):

\[\sum_{n =1}^{\infty} \frac{ (-2)^{n} }{ e^{n+1} } = \dfrac{1}{e}\sum_{n =1}^{\infty} \frac{ (-2)^n}{ e^n }=\dfrac{1}{e}\sum_{n =1}^{\infty} \left(\dfrac{-2}{e}\right)^{n} \]Yes?

OpenStudy (anonymous):

How would I go from there?

geerky42 (geerky42):

Familiar with formula for sum of an infinite geometric series?

OpenStudy (phi):

e > 2 so you have a geometric series with |r| < 1

OpenStudy (anonymous):

Oh... I see. Thanks!

geerky42 (geerky42):

Be careful; the formula is \[\sum_{n=0}^\infty r^n = \dfrac{1}{1-r}\] Here, you have bottom limit of summation \(n=1\) instead of \(n=0\).

geerky42 (geerky42):

And welcome.